# SPEED: Scalable, Precise, and Efficient Concept Erasure for Diffusion Models

### v2

@@ C.2 Erasure Configurations
Appendix D Additional Experiments
D.1 More Demonstrations
We also evaluate the CLIP score (CS) before and after concept erasure to assess the erasure efficacy. As shown in Fig. 8, our method achieves successful erasure of specific concepts such as Snoopy and Mickey while maintaining moderate CS values (24.18 and 23.44, respectively). This indicates that effective erasure does not require minimizing CS to an extreme. In contrast, RECE obtains the lowest CS (19.79 and 18.75), but this is achieved at the cost of overly aggressive erasure. For example, transforming Snoopy into an unrecognizable image and replacing Mickey with a generic human figure. While such strategies may enhance erasure efficacy, they also risk compromising prior knowledge. This trade-off is reflected in higher non-target FIDs, as shown in Tables 1 and 2.
D.2 Complete Results on Few-Concept Erasure
Figure 8: Comparison of CLIP scores (CS) across different erasure methods. We compare the results in erasing Snoopy and Mickey, and highlight the corresponding CS in blue. Our method achieves successful concept erasure with moderate CS values. In contrast, RECE achieves the lowest CS by enabling more aggressive erasure. For example, removing Snoopy to the extent of producing a semantically void image, and changing Mickey into a generic person. We argue that such over-erasure unnecessarily compromises prior preservation as evidenced by Tables 1 and 2.
@@ D.3 Comparison on More Baselines
The quantitative comparative results are shown in Tables 5 and 6, where our method consistently achieves the best prior preservation compared to all compared baselines. Even equipped with Facilitated Transport (i.e., SPM w/ FT), our method achieves the lowest non-target FID (e.g., on Pikachu and Hello Kitty). This superiority amplifies as the number of target concepts increases as shown in Table 7. For example, with the number of target concepts increasing from 1 to 3, our FID in Pikachu rises from 16.81 to 21.40 (4.59 ↑), while SPM increases from 19.82 to 37.51 (17.69 ↑), where a similar pattern is observed in Hello Kitty (Our 4.48 ↑ v.s. SPM’s 15.68 ↑). Once removing the Facilitated Transport module, SPM w/o FT shows poorer prior preservation with rapidly increasing FIDs (highlighted in red in Table 5). More qualitative results are shown in Fig. 9.
@@ D.5 Ablation Studies
Appendix E Limitation
Despite the promising results, our work still has limitations. SPEED relies on linear null-space projections, which may not fully capture the nonlinear interactions between concepts in large diffusion models. In practice, this can lead to imperfect preservation when erasing highly entangled or stylistically subtle concepts. In addition, our evaluation mainly covers benchmarks with explicit or moderately implicit concepts; the effectiveness on more abstract (e.g., freedom), compositional (e.g., a blue cat), or culturally nuanced (e.g., Día de los Muertos) concepts remains less explored. Finally, although our method scales efficiently to 100 concepts, extending it to even larger-scale or continual erasure settings may require additional mechanisms.

### v4

@@ D.3 Comparison on More Baselines
The quantitative comparative results are shown in Tables 6 and 7, where our method consistently achieves the best prior preservation compared to all compared baselines. Even equipped with Facilitated Transport (i.e., SPM w/ FT), our method achieves the lowest non-target FID (e.g., on Pikachu and Hello Kitty). This superiority amplifies as the number of target concepts increases as shown in Table 8. For example, with the number of target concepts increasing from 1 to 3, our FID in Pikachu rises from 16.81 to 21.40 (4.59 ↑), while SPM increases from 19.82 to 37.51 (17.69 ↑), where a similar pattern is observed in Hello Kitty (Our 4.48 ↑ v.s. SPM’s 15.68 ↑). Once removing the Facilitated Transport module, SPM w/o FT shows poorer prior preservation with rapidly increasing FIDs (highlighted in red in Table 6). More qualitative results are shown in Fig. 9.
@@ D.5 Ablation Studies
Appendix E LLM Usage Statement
We use large language models (LLMs) as an auxiliary tool during preparing this work. The LLM is employed to generate artistic style templates for the artistic style erasure task (see Appx. C.1) and to refine the clarity and readability of certain parts of the manuscript, such as polishing grammar, improving fluency, and standardizing terminology. In addition, LLMs are occasionally used to suggest alternative phrasings when writing sections like the introduction and related work, but the final narrative, arguments, and presentation choices are made solely by the authors. All methodological ideas, theoretical derivations, experiment designs, and analyses are developed independently by the authors without assistance from LLMs. We do not rely on LLMs for generating novel research ideas, conducting experiments, interpreting results, or writing technical content. The role of LLMs is purely supportive and limited to stylistic refinement and auxiliary text generation, and thus they are not regarded as scientific contributors to this paper.
Appendix F Limitation
Despite the promising results, SPEED is designed with linear null-space projections, which may not fully capture the nonlinear interactions between concepts in large diffusion models. In practice, this can lead to imperfect preservation when erasing highly entangled or stylistically subtle concepts. In addition, our evaluation mainly covers benchmarks with explicit or implicit concepts; the effectiveness on more abstract (e.g., freedom), compositional (e.g., a blue cat), or cultural (e.g., Día de los Muertos) concepts remains less explored. Finally, although our method scales efficiently to 100 concepts, extending it to even larger-scale or continual erasure may require additional mechanisms.